# Revealing Superconducting Chiral Edge Modes via Resistance Distributions

### v1

@@ Abstract
Inducing superconducting correlations in quantum anomalous Hall (QAH) states offers a promising route to realize topological superconductivity with chiral Majorana edge modes. However, the definitive identification of these modes is challenging. Here we propose detecting superconducting chiral edge modes via the probability distribution of the resistance, or equivalently the charge transmission of QAH-superconductor heterojunctions. Remarkably, the distribution for coherent edge exhibits distinct characteristics for different topological superconducting phases in sufficiently long junctions, and this difference remains robust against weak decoherence. These findings provide new insights into transport phenomena beyond the clean limit and highlight the resistance distribution as a compelling signature for distinguishing topological superconducting phases.
@@
The superconducting chiral edge modes, which are unidirectional and dissipationless Bogoliubov edge states, could emerge along the boundaries of a two-dimensional chiral topological superconductor (TSC) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and have potential applications in quantum information [11, 12, 13, 14, 15, 16]. A promising approach to realizing these modes is to induce superconducting pair correlation in the chiral edge states of quantum Hall (QH) or quantum anomalous Hall (QAH) insulators via the proximity effect [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Distinctive transport phenomena have been observed in the QH-SC and QAH-SC junctions [24, 25, 27, 26, 28] due to the charge-neutral nature of these edge modes [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. A significant goal is to achieve $p+ip$ TSC, which hosts a single chiral Majorana edge mode exhibiting quantized transport and interference signatures [34, 35, 36, 37, 52]. However, its experimental confirmation remains elusive. At the same time, disorder and decoherence are inherent in realistic experimental devices [28, 26], making it essential to study transport signatures beyond the idealized clean limit. In particular, understanding the resistance distribution under these effects is vital [43, 48, 49], especially as such distributions can now be measured directly in experiments [26]. Theoretically, an intriguing question is whether these distributions reveal novel behaviors due to the presence of exotic Majorana edge states. Experimentally, it is crucial to determine whether the distribution, beyond just its mean value, provides stronger evidence to distinguish between different topological phases in SC.
@@
In this Letter, we present a comprehensive analytical investigation of disorder-induced resistance distributions in QAH-SC junctions, supported by extensive numerical simulations. Using the transfer matrix formalism, we show that the quantum transport of edge modes under disorder can be intuitively described as random rotations on the Bloch sphere. This framework offers a more rigorous approach than that of [43], which relied on simplifications from renormalization group analysis. We specifically consider QAH-SC heterojunctions commonly implemented in experiments as illustrated in Fig. 1, where the QAH insulator has a Chern number $C=1$, hosting a single chiral electron edge mode [53, 54, 55, 56, 57], while the SC region is characterized by a Bogoliubov-de Gennes (BdG) Chern number $N=1$ or $N=2$, corresponding to transport mediated by one or two chiral Bogoliubov edge modes (CBEM), respectively. The key results are summarized in Table 1, which outlines the probability distributions of charge transmission (defined in Eq. (1)) across different scenarios. Notably, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEM exhibit distinct characteristics. Moreover, these differences remain robust against decoherence, suggesting that resistance distributions serve as a compelling signature to distinguish between different topological phases of SC.
@@
Table 1: Summary of the probability distributions of charge transmission fraction $T$ in two commonly implemented heterojunction configurations. $f(T)$: the probability density function of $T$. $\mathcal{N}$: normal distribution. $\chi^{2}$: chi-squared distribution. $U[-1,1]$: uniform distribution over $[-1,1]$.
@@
Edge transport with $N=2$ CBEM. By stacking a normal SC onto the middle region of a $C=1$ QAH insulator, one can experimentally implement a QAH-SC-QAH heterojunction in Fig. 1(a). We first consider the middle region of the heterojunction to be in the same topological phase as the $C=1$ QAH state with broken charge $U(1)$ symmetry, namely a TSC with BdG Chern number $N=2$, which would be the case when the SC proximity gap is smaller than the insulating QAH bulk gap [6, 37]. In Nambu basis $(c_{k}^{\dagger},~{}c_{-k})^{\text{T}}$, the effective edge Hamiltonian for TSC is described as $H_{\text{edge}}=vk\tau_{0}-\mu\tau_{z}+\Delta\tau_{x}$, where $v$ is the edge mode velocity, $\mu$ is the chemical potential, and $\Delta=\Delta_{0}k$ is the SC pairing amplitude. $\tau_{x,y,z}$ are Pauli matrices in Nambu space. Accordingly, the single chiral electron edge mode of the QAH state splits into a particle-hole conjugate pair of CBEM of the TSC state, also knows as chiral Andreev edge modes [24]. In Fig. 1(a), normal transmission and Andreev transmission [58] refer to the processes where an electron from the QAH region enters the SC and exits as either an electron or a hole, with transmission coefficients denoted by $T_{\text{ee}}$ and $T_{\text{eh}}$, respectively. We define the charge transmission fraction [59] from lead 2 to 3 as,
@@
The presence of static disorder induces spatial variations in the potential and pairing amplitude experienced by the edge modes. This effect is effectively described by random fluctuations of the parameters $(v,\mu,\Delta_{0})$ in $H_{\text{edge}}$, making $T$ a random variable governed by a probability distribution—this forms the central focus of our study. Experimentally, this distribution can be observed by varying the magnetic field or the electron density [60, 48, 26]. We divide the SC region into $L$ segments labeled by the index $\ell=1,2,...,L$, where $L\propto\mathcal{L}/\xi_{d}$. Here $\mathcal{L}$ denotes the junction length and $\xi_{d}$ is the disorder correlation length. Within each segment, the parameters $(v,\mu,\Delta_{0})$ remain constant in the local Hamiltonian $H_{\ell}$, as depicted in Fig. 1(b). The final state is obtained via the transfer matrix method [33, 61, 62]
$\Psi_{f}=(\psi_{\text{e}},\psi_{\text{h}})^{\text{T}}=\mathbb{T}_{L}\cdots\mathbb{T}_{2}\mathbb{T}_{1}\Psi_{i}\equiv\mathbb{T}\Psi_{i},$ (2)
where $\Psi_{i}=(1,0)^{\text{T}}$ is the incident state. The transfer matrix for the $\ell$-th SC segment $\mathbb{T}_{\ell}$ is given by $\mathbb{T}_{\ell}=P_{\ell}^{-1}\Lambda_{\ell}P_{\ell}$, where $\Lambda_{\ell}$ represents the propagation matrix of the segment and $P_{\ell}$ transforms the Nambu basis into the eigenbasis of $H_{\ell}$. $\mathbb{T}_{\ell}$ is intuitively interpreted as a rotation on the Bloch sphere, with a rotation angle $\alpha_{\ell}$ around an axis $\mathbf{n}_{\ell}$, as depicted in Fig. 1(c). Here we consider $L>2\ell_{d}/a$, where $\ell_{d}$ is the decay length of CBEM wavefuntion, and $a$ is the lattice constant. Therefore, the crossed Andreev reflection and elastic co-tunneling processes are neglected [63, 46, 64, 65]. As detailed in [62], both $\alpha_{\ell}$ and $\mathbf{n}_{\ell}$ exhibit small fluctuations, with standard deviations satisfying $\delta\alpha_{\ell}\ll 2\pi$ and $|\delta\mathbf{n}_{\ell}|\ll 1$. The cumulative effect of these random rotations can be approximated as a single rotation by an angle $\alpha$ around an axis $\mathbf{n}\equiv(n_{x},n_{y},n_{z})$, such that $\mathbb{T}\propto e^{i\alpha(\mathbf{n}\cdot\bm{\sigma})/2}$, where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are Pauli matrices. The final state then takes the form $\Psi_{f}=(\cos{(\alpha/2)}+in_{z}\sin{(\alpha/2)},i(n_{x}+in_{y})\sin{(\alpha/2)})^{\text{T}}$, and the charge transmission fraction is given by $T=|\psi_{\text{e}}|^{2}-|\psi_{\text{h}}|^{2}=n_{z}^{2}+(1-n_{z}^{2})\cos{\alpha}$.
@@
Figure 1: (a) The QAH-SC-QAH junction with $N=2$ CBEM. The orange and dashed blue arrows represent the chiral electron edge mode of QAH state and the CBEM of TSC state, respectively. (b) Transfer matrix construction for the junction shown in (a). (c) Evolution of the state on the Bloch sphere. Both the axis and rotation angle exhibit weak fluctuations. The north (south) pole on the Bloch sphere represents a pure electron (hole) state. (d,e) The QAH-SC-QAH junction with $N=1$ CBEM, and the corresponding transfer matrix construction. (f) The QAH-SC junction with $N=1$ CBEM.
The distribution of $T$ exhibits distinct yet general behaviors depending on the length of SC region, as analyzed in the following cases.
Case (i): For a relatively short QAH-SC-QAH junction (small $L$ as $2\ell_{d}/a<L<10^{2}$) in Fig. 1(a), $T$ approximately follows either a normal distribution ($\mathcal{N}$), or a chi-squared distribution ($\chi^{2}$).
In this regime, fluctuations in the rotation axis are negligible, allowing us to approximate $\mathbf{n}\approx\overline{\mathbf{n}}_{\ell}$. As a result, the state undergoes sequential rotations around a nearly fixed axis, with each segment contributing a random rotation angle. The total rotation angle is $\alpha=\sum_{\ell=1}^{L}\alpha_{\ell}$. By the central limit theorem, after multiple segments, $\alpha$ follows a normal distribution with mean $\overline{\alpha}=L\overline{\alpha}_{\ell}$ and standard deviation $\delta{\alpha}=\sqrt{L}\delta{\alpha}_{\ell}$, irrespective of the specific distribution of the individual rotation angles $\alpha_{\ell}$. When $L$, and thus $\delta\alpha$, is small, we can expand $\alpha$ around $\overline{\alpha}$ up to second order: $\cos{\alpha}\approx\cos{\overline{\alpha}}-\sin{\overline{\alpha}}\Delta\alpha-\cos{\overline{\alpha}}(\Delta\alpha)^{2}/2$, where $\Delta\alpha$ follows a normal distribution. Therefore, $T$ approximately follows a normal distribution when $\sin{\overline{\alpha}}$ is not close to zero, and a chi-squared distribution when $\sin{\overline{\alpha}}\approx 0$ [62]. As the junction length varies, the distribution of $T$ oscillates between normal and chi-squared distributions. This behavior extends beyond the results in [48], where only a normal distribution was proposed.
Case (ii): For a long QAH-SC-QAH junction (large $L$ as $L>10^{4}$) in Fig. 1(a), $T$ approximates a continous uniform distribution over $[-1,1]$, denoted as $U[-1,1]$.
As $L$ increases, the fluctuations of the rotation axis become significant, causing the state to spread across the entire Bloch sphere, as illustrated in Fig. 1(c). The ergodicity theorem [66] states that a point undergoing repeated rotations about two noncollinear axes eventually becomes uniformly distributed over the sphere. Since the fluctuating rotation axes in our system satisfy this condition, the final state $\Psi_{f}$ becomes equidistributed on the Bloch sphere, leading to $T\sim U[-1,1]$. A more rigorous derivation is provided in [62].

### v2

@@ Abstract
Inducing superconducting correlations in quantum anomalous Hall (QAH) states offers a promising route to realize topological superconductivity with chiral Majorana edge modes. However, the definitive identification of these modes is challenging. Here we propose detecting superconducting chiral edge modes via the probability distribution of the resistance, or equivalently the charge transmission of QAH-superconductor heterojunctions. Remarkably, the distribution for coherent edge exhibits distinct characteristics for different topological superconducting phases in sufficiently long junctions, and this difference remains robust against weak decoherence. These findings provide insights into transport phenomena beyond the clean limit and highlight the resistance distribution as a compelling signature for distinguishing topological superconducting phases.
I Introduction
@@
The superconducting chiral edge modes, which are unidirectional and dissipationless Bogoliubov edge states, could emerge along the boundaries of a two-dimensional chiral topological superconductor (TSC) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and have potential applications in quantum information [11, 12, 13, 14, 15, 16]. A promising approach to realizing these modes is to induce superconducting pair correlation in the chiral edge states of quantum Hall (QH) or quantum anomalous Hall (QAH) insulators via the proximity effect [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Distinctive transport phenomena have been observed in the QH-SC and QAH-SC junctions [24, 25, 27, 26, 28] due to the charge-neutral nature of these edge modes [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. A significant goal is to achieve $p+ip$ TSC, which hosts a single chiral Majorana edge mode exhibiting quantized transport and interference signatures [35, 36, 37, 38, 53]. However, its experimental confirmation remains elusive. At the same time, disorder and decoherence are inherent in realistic experimental devices [28, 26], making it essential to study transport signatures beyond the idealized clean limit. In particular, understanding the resistance distribution under these effects is vital [44, 49, 50], especially as such distributions can now be measured directly in experiments [26]. Theoretically, an intriguing question is whether these distributions reveal novel behaviors due to the presence of exotic Majorana edge states. Experimentally, it is crucial to determine whether the distribution, beyond just its mean value, provides stronger evidence to distinguish between different topological phases in SC.
@@
Here, we present a comprehensive analytical investigation of disorder-induced resistance distributions in QAH-SC junctions, supported by extensive numerical simulations. Using the transfer matrix formalism, we show that the quantum transport of edge modes under disorder can be intuitively described as random rotations on the Bloch sphere. This framework offers a more rigorous approach than that of Ref. [44], which relied on simplifications from renormalization group analysis. We specifically consider QAH-SC heterojunctions commonly implemented in experiments as illustrated in Fig. 1, where the QAH insulator has a Chern number $C=1$, hosting a single chiral electron edge mode [54, 55, 56, 57, 58], while the SC region is characterized by a Bogoliubov-de Gennes (BdG) Chern number $N=1$ or $N=2$, corresponding to transport mediated by one or two chiral Bogoliubov edge modes (CBEMs), respectively. The key results are summarized in Table 1, which outlines the probability distributions of charge transmission [defined in Eq. (2)] across different scenarios. Notably, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEMs exhibit distinct characteristics. Moreover, these differences remain robust against decoherence, suggesting that resistance distributions serve as a compelling signature to distinguish between different topological phases of SC.
@@
Table 1: Summary of the probability distributions of charge transmission fraction $T$ in two commonly implemented heterojunction configurations. $f(T)$: the probability density function of $T$; $\mathcal{N}$: normal distribution; $\chi^{2}$: chi-squared distribution; and U[1,1]U[-1,1]: uniform distribution over [1,1][-1,1].
@@
The paper is organized as follows. We use transfer matrix formalism to solve the superconducting edge-mode transport problem of both $N=1$ and $N=2$ CBEM cases for different junction configurations in Sec. II. Section III presents the derivation of the distribution function of the corresponding charge transmission fraction for each scenario. We use two-dimensional lattice models to numerically simulate this transport problem and verify our analytical results in Sec. IV. In Sec. V, the effect of decoherence and particle loss on the distribution is studied, and we find that the qualitative conclusion in Sec. III remain robust against weak decoherence. Finally, we conclude with some further discussions in Sec. VI.
II Transport theory of superconducting edge modes
II.1 Edge transport with $N=2$ CBEMs
By stacking a normal SC onto the middle region of a $C=1$ QAH insulator, one can experimentally implement a QAH-SC-QAH heterojunction in Fig. 1(a). We first consider the middle region of the heterojunction to be in the same topological phase as the $C=1$ QAH state with broken charge $U(1)$ symmetry, namely a TSC with BdG Chern number $N=2$, which would be the case when the SC proximity gap is smaller than the insulating QAH bulk gap [6, 38]. In Nambu basis $(\psi(x),~\psi^{\dagger}(x))^{\text{T}}$, the effective edge Hamiltonian is described as [59]
$H(x)=v(x)k\tau_{0}-\mu(x)\tau_{z}+\Delta(x)k\tau_{x},$ (1)
where $k\equiv-i\partial_{x}$, $v(x)$ is the edge-mode velocity, $\mu(x)$ is the chemical potential, and $\Delta(x)k$ is the SC pairing amplitude, which is finite in the TSC regime and vanishes in the QAH regime. Here, $\tau_{x,y,z}$ denote the Pauli matrices in Nambu space and $\tau_{0}$ is the identity matrix. Accordingly, the single chiral electron edge mode of the QAH state splits into a particle-hole conjugate pair of CBEM of the TSC state, also known as chiral Andreev edge modes [24]. In Fig. 1(a), normal transmission and Andreev transmission [60] refer to the processes in which an electron from the QAH region enters the SC and exits as an electron or a hole, with transmission coefficients denoted by $T_{\text{ee}}$ and $T_{\text{eh}}$, respectively. We define the charge transmission fraction [61] from lead 2 to 3 as,
@@ II.1 Edge transport with $N=2$ CBEMs
Figure 1: (a) The QAH-SC-QAH junction with $N=2$ CBEM. The orange and dashed blue arrows represent the chiral electron edge mode of QAH state and the CBEM of TSC state, respectively. (b) Transfer matrix construction for the junction shown in (a). (c) Evolution of the state on the Bloch sphere. Both the axis and rotation angle exhibit weak fluctuations. The north (south) pole on the Bloch sphere represents a pure electron (hole) state. (d), (e) The QAH-SC-QAH junction with $N=1$ CBEM, and the corresponding transfer matrix construction. (f) The QAH-SC junction with $N=1$ CBEM.
The presence of static disorder induces spatial variations in the potential and pairing amplitude experienced by the edge modes. This effect is effectively described by random fluctuations of the parameters $(v,\mu,\Delta_{0})$ in $H_{\text{edge}}$, making $T$ a random variable governed by a probability distribution—this forms the central focus of our study. Experimentally, this distribution can be observed by varying the magnetic field or the electron density [62, 49, 26]. $T$ can be obtained via the transfer matrix method [34, 63]: We start from the eigenvalue equation for edge-mode transport: $H(x)\Psi(x)=\varepsilon\Psi(x)$, where $\varepsilon$ is the energy of the incident electron. By defining the following operators:
$J(x)=\frac{\partial H}{\partial k}=\begin{pmatrix}v(x)&\Delta(x)\\ \Delta(x)&v(x)\end{pmatrix}$, (3)
$J^{\frac{1}{2}}(x)=\begin{pmatrix}c_{0}(x)&c_{x}(x)\\ c_{x}(x)&c_{0}(x)\end{pmatrix},~\text{where}$
$c_{0,x}(x)=\frac{\sqrt{v(x)+\Delta(x)}\pm\sqrt{v(x)-\Delta(x)}}{2},$
which satisfies $\left[J^{\frac{1}{2}}(x)\right]^{2}=J(x)$, the TSC effective edge Hamiltonian can be diagonalized as
$\tilde{H}(x)$ $=J^{-\frac{1}{2}}(x)H(x)J^{-\frac{1}{2}}(x)$ (4)
$=k\tau_{0}-\frac{\mu(x)}{\sqrt{[v(x)]^{2}-[\Delta(x)]^{2}}}\tau_{z}.$
Here, $J^{-\frac{1}{2}}(x)$ denotes the inverse of $J^{\frac{1}{2}}(x)$. Transforming the eigenvalue equation to $\tilde{H}(x)\tilde{\Psi}(x)=\varepsilon\tilde{\Psi}(x)$ with $\tilde{\Psi}(x)=J^{\frac{1}{2}}(x)\Psi(x)$, one can straightforwardly solve it to obtain
$\Psi_{f}=J^{-\frac{1}{2}}(x_{f})e^{\int_{x_{i}}^{x_{f}}A(x)dx}J^{-\frac{1}{2}}(x_{i})\Psi_{i}=e^{\int_{x_{i}}^{x_{f}}A(x)dx}\Psi_{i},$ (5)
$A(x)=i\left[\varepsilon J^{-1}(x)+\frac{1}{[v(x)]^{2}-[\mu(x)]^{2}}\right]\mu(x)\tau_{z},$
where $x_{i(f)}$ are the initial (final) positions of the TSC region, $\Psi_{f}=\Psi(x_{f})=(\psi_{\text{e}},\psi_{\text{h}})^{\text{T}}$ is the final state, and $\Psi_{i}=\Psi(x_{i})=(1,0)^{\text{T}}$ is the incident state. In the first line we have used $J^{-\frac{1}{2}}(x_{i}-0^{+})=J^{-\frac{1}{2}}(x_{f}+0^{+})$ for the QAH regime, both proportional to $\tau_{0}$. Disorder in the TSC region induces fluctuations in $v(x)$, $\mu(x)$, and $\Delta(x)$, which render the integral in Eq. (5) analytically intractable. For simplicity, we divide the SC region into $L$ segments labeled by the index $\ell=1,2,...,L$, where $L\propto\mathcal{L}/\xi_{d}$. Here $\mathcal{L}$ denotes the junction length and $\xi_{d}$ is the segment length. Within each segment, the parameters $(v_{\ell},\mu_{\ell},\Delta_{\ell})$ remain constant in the local Hamiltonian $H_{\ell}$, as depicted in Fig. 1(b). The integral in Eq. (5) is then converted into a sum, yielding the final state:
$\Psi_{f}=e^{A_{L}}\cdots e^{A_{2}}e^{A_{1}}\Psi_{i}=P_{L}^{-1}\Lambda_{L}P_{L}\cdots P_{1}^{-1}\Lambda_{1}P_{1}\Psi_{i},$ (6)
In the second equation, the $A_{\ell}$ matrices are diagonalized piecewise. Denoting the transfer matrix as $\mathbb{T}_{\ell}=P_{\ell}^{-1}\Lambda_{\ell}P_{\ell}$, we obtain
$\Psi_{f}=\mathbb{T}_{L}\cdots\mathbb{T}_{2}\mathbb{T}_{1}\Psi_{i}\equiv\mathbb{T}\Psi_{i},$ (7)
where $\Lambda_{\ell}$ represents the propagation matrix of each segment and $P_{\ell}$ transforms the Nambu basis into the eigenbasis of $H_{\ell}$.
The explicit expressions for $P_{\ell}$ and $\Lambda_{\ell}$ are
$P_{\ell}$ $=\begin{pmatrix}\frac{\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}-\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}}{\sqrt{\Delta_{\ell}^{2}\varepsilon^{2}+\left(\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}-\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}\right)^{2}}}&-\frac{\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}+\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}}{\sqrt{\Delta_{\ell}^{2}\varepsilon^{2}+\left(\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}+\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}\right)^{2}}}\\ \frac{\Delta_{\ell}\varepsilon}{\sqrt{\Delta_{\ell}^{2}\varepsilon^{2}+\left(\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}-\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}\right)^{2}}}&\frac{\Delta_{\ell}\varepsilon}{\sqrt{\Delta_{\ell}^{2}\varepsilon^{2}+\left(\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}+\mu_{\ell}\sqrt{v_{\ell}^{2}-\Delta_{\ell}^{2}}\right)^{2}}}\end{pmatrix}$, (8)
$\Lambda_{\ell}$ $=\begin{pmatrix}e^{i\frac{v_{\ell}\varepsilon-\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}}{v_{\ell}^{2}-\Delta_{\ell}^{2}}L_{\ell}}&0\\ 0&e^{i\frac{v_{\ell}\varepsilon+\sqrt{(v_{\ell}^{2}-\Delta_{\ell}^{2})\mu_{\ell}^{2}+\Delta_{\ell}^{2}\varepsilon^{2}}}{v_{\ell}^{2}-\Delta_{\ell}^{2}}L_{\ell}}\end{pmatrix}$,
[…]
Applying a gauge transformation to make $\psi_{\text{e}}$ real, the wavefunction can be parametrized as $(\cos(\theta/2),~e^{i\phi}\sin(\theta/2))^{\text{T}}$, where $\theta$ and $\phi$ are the polar and azimuthal angles on the Bloch sphere, respectively. In this representation, $\Psi_{i}$ lies at the north pole. The $2\times 2$ unitary matrix $\mathbb{T}_{\ell}$ can be parametrized as $N_{\ell}e^{i\alpha_{\ell}(\mathbf{n}_{\ell}\cdot\bm{\sigma})/2}$, where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$. This indicates that $\mathbb{T}_{\ell}$ corresponds to a rotation on the Bloch sphere by an angle $\alpha_{\ell}$ around the axis $\mathbf{n}_{\ell}$, as shown in Fig. 1(c). From Eq. (8), we find
$N_{\ell}$ $=e^{\frac{iL_{\ell}v_{\ell}\varepsilon}{v_{\ell}^{2}-\Delta_{\ell}^{2}}},\quad\alpha_{\ell}=\frac{2L_{\ell}\Delta_{\ell}\varepsilon\sqrt{1+s_{\ell}}}{v_{\ell}^{2}-\Delta_{\ell}^{2}}~\text{mod}~2\pi,$ (9)
$\mathbf{n}_{\ell}$ $=\left(\frac{1}{\sqrt{1+s_{\ell}}},~0,~\frac{1}{\sqrt{1+s_{\ell}^{-1}}}\right)^{\text{T}},$
where $s_{\ell}\equiv\frac{\mu_{\ell}^{2}}{\varepsilon^{2}}\left(\frac{v_{\ell}^{2}}{\Delta_{\ell}^{2}}-1\right)$. The rotation axis always lies in the xozxoz plane. When the parameters of the effective Hamiltonian fluctuate slightly due to disorder, both $\alpha_{\ell}$ and $\mathbf{n}_{\ell}$ exhibit small variations, with standard deviations satisfying $\delta\alpha_{\ell}\ll 2\pi$ and $|\delta\mathbf{n}_{\ell}|\ll 1$, particularly when $s_{\ell}$ is large [64]. Hence, the total transport process can be regarded as the evolution of a point at the north pole under successive rotations. The cumulative effect of these random rotations can be represented by a single rotation by an angle $\alpha$ around an axis $\mathbf{n}\equiv(n_{x},n_{y},n_{z})$, such that $\mathbb{T}\propto e^{i\alpha(\mathbf{n}\cdot\bm{\sigma})/2}$. The final state then takes the form $\Psi_{f}=(\cos{(\alpha/2)}+in_{z}\sin{(\alpha/2)},i(n_{x}+in_{y})\sin{(\alpha/2)})^{\text{T}}$, and the charge transmission fraction is given by
$T=|\psi_{\text{e}}|^{2}-|\psi_{\text{h}}|^{2}=n_{z}^{2}+(1-n_{z}^{2})\cos{\alpha}.$ (10)
Although the resulting rotation is generally complicated, one simple limit can be identified: as $\varepsilon\to 0$, $s_{\ell}\to\infty$ and $\mathbf{n}_{\ell}$ aligns with the $z$ axis. Since the initial state lies at the north pole, the final state also remains at the north pole, leading to no Andreev conversion and $T=1$. This behavior arises from the Pauli exclusion principle when forming the Cooper pair [65, 66, 34, 63]. Furthermore, this Bloch sphere picture shows that $T$ depends on the incident energy $\varepsilon$. Larger incident energy may increase the probability of Andreev conversion. In the following, we therefore restrict to finite $\varepsilon$ to ensure Andreev conversion.
II.2 Edge transport with $N=1$ CBEM